# ShoppingComp: Are LLMs Really Ready for Your Shopping Cart?

### v1

@@ 4 Evaluations
We organize our evaluation into four parts. First, we introduce grading metrics, including Answer Match for product retrieval, rubric-based measures for report quality and safety rubric pass rate. Next, we detail the models and settings, covering both LLMs and commercial DeepResearch products under unified evaluation protocols. We then present performance results, highlighting product accuracy, report quality, and safety-critical scenarios. Finally, we provide analysis, examining models’ product searching ability, report generation and robustness in safety-critical traps.
@@ 4.1 Grading
For a given question $q\in Q$, we first decompose it into a set of distinct scenarios $S$. Alongside this, we establish a collection of reasoned and validated rubrics, denoted as $R$ Subsequently, we identify a set of satisfying products, $P$ For each rubric $r_{i}\in R$ and each product $p_{j}\in P$, we provide the corresponding annotator-verified evidence, $e_{ij}\in E$. Each piece of evidence $e_{ij}\in E$ can exist in various forms. Consequently, the complete evidence set $E$ is multi-modal, consisting of text, URLs, and images. All the following metrics are judged using LLM-as-a-Judge framework (Sec. 3.3) to ensure semantic alignment and consistency.
@@ 4.1.1 Browse Products Score
Given a question $q$, a set of rubrics $R$, and a set of ground-truth products $P$, we use answer match (AM) to calculate the model’s predicted product set, denoted as $\hat{P}$. This method assesses the semantic correspondence between the predicted products and the ground-truth products. We leverage LM-as-a-Judge to determine if a predicted product $\hat{P}$ semantically matches any product in the ground-truth set $P$. Since our ground-truth data consists solely of correct products, we evaluate performance using the standard metrics of Precision, Recall, and the F1-score.
@@ 4.1.2 Report Score
To systematically and comprehensively evaluate the quality of product recommendation report, three dimensions are considered: 1) Scenario Coverage, 2) Selection Accuracy, and 3) Rationale Validity.
Scenario Coverage. This metric measures how well the report identifies the user’s demand scenarios. Let the set of ground-truth scenarios be $S$ and the set of scenarios predicted in the report be $\hat{S}$. We use LLM-as-a-Judge to determine the semantic alignment indicating if predicted scenario $\hat{s}_{n}\in\hat{S}$ correctly matches the ground-truth scenario $s_{m}\in S$. Then, we calculate the standard metrics of Precision, Recall, and the F1-score. This score evaluates the comprehensiveness and accuracy of the model’s demand understanding and instruct following.
Selection Accuracy. This metric evaluates the quality of the products recommended in the report. It measures the proportion of recommended products that satisfy the user’s requirements, as defined by a set of $N$ rubrics. Let the list of model-recommended products be $\hat{P}=\{\hat{p}_{1},\dots,\hat{p}_{J}\}$. LLM-as-a-Judge provides a correctness judgment, $c_{nj}$, which is 1 if product $\hat{p}_{j}$ satisfies rubric $r_{n}$ and 0 otherwise. The overall quality of the product list is measured by the Satisfaction of Products (SoP):
A higher SoP score indicates that the report recommends more suitable products.
Rationale Validity. This metric assesses whether the reasoning provided in the report is correct and logically sound. For each requirement rubric $r_{n}$ and a ground-truth product $p_{j}$, an LLM-as-a-Judge evaluates the corresponding reasoning in the report. The judge outputs a boolean score, $v_{nj}$, which is 1 if the reasoning is valid and 0 otherwise. The overall Rationale Validity (RV) is the average score across all $N$ rubrics and $L$ ground-truth products:
$\text{RV}=\frac{\sum_{n=1}^{N}\sum_{j=1}^{L}v_{nj}}{N\cdot L}$ (2)
This metric directly measures the factuality and logical integrity of the report’s explanations.
@@ 4.1.3 Safety Rubric Pass Rate
The Safety Rubric Pass Rate calculates the percentage of “safety trap” questions that are successfully addressed by the model. $|Q_{s}|$ is the total number of safety trap questions $Q_{s}$, $R_{i}$ is the set of all required ground-truth safety rubrics for that question. The evaluation for each question is measured by LLM verifier.
$\text{PassRate}_{\text{Safety}}=\frac{\sum_{q_{i}\in Q_{s}}\text{LLM}(R_{i},\hat{\text{Report}}_{i})}{|Q_{s}|}$ (3)
4.2 Evaluation of models
4.2.1 Models and Settings
Models: We evaluate a broad spectrum of both open-source and proprietary offerings on ShoppingComp to offer a diverse and comprehensive benchmark, distinguishing between foundational models accessed via API and deep research products. Our evaluation includes a range of language model APIs:  gpt-5-2025-08-07 (OpenAI, 2025b),  gpt-4o-2024-11-20 (OpenAI, 2024),  gemini-2.5-pro (Google, 2025b),  gemini-2.5-flash (Google, 2025a),  deepseek-V3.1-0821 (DeepSeek, 2025),  claude4-sonnet (Anthropic, 2025b), and  claude4-opus (Anthropic, 2025a). For these models, we treat them as black-box services, applying identical prompt templates and evaluation protocols for a fair comparison. In contrast, for deep research products like ChatGPT DeepResearch (OpenAI, 2025a) and Gemini DeepResearch (Google, 2025c), we assess their end-to-end performance as holistic systems, focusing on the final output provided to the user.
Experimental Settings: We use three representative LLM tools: search, powered by SerpAPI [^1] https://serpapi.com/ for Google Search access, as well as link_reader and link_summary, which leverage Firecrawl [^2] https://www.firecrawl.dev/ to retrieve the full content and a concise summary of a webpage, respectively. Each LLM is allowed to invoke these tools as external function calls during the reasoning process, following the paradigm of tool-augmented language modeling (Schick et al., 2023; Patil et al., 2023; Luo et al., 2023). For a fair comparison, we conduct five independent runs per test case, reporting the average performance across runs. We emphasize the use of averages rather than best-of-N results, as average performance more faithfully reflects real-world deployment scenarios, where consistency and robustness are often more critical than peak outcomes.
4.2.2 Performance
Performance of products and report score
A summary of the overall results on product retrieval and report quality is given in Tab. 2. Human experts achieve an Answer Match F1 of 25.7% and a Report Validity (RV) of 90.9%. In contrast, annotators achieve similar retrieval accuracy but slightly lower RV (85.2%). Among LLMs, GPT-5 substantially outperforms all other models with 11.2% F1 on product retrieval and 90.3% RV, demonstrating strong reasoning and factuality in report explanations. However, even GPT-5 falls short of human-level retrieval, highlighting the challenge of translating vague consumer needs into precise search constraints. Other models, such as Gemini-2.5-pro and Claude-4-opus, achieve moderate RV scores (83.6% and 77.3% respectively) but suffer from low retrieval recall. DeepResearch products deliver significantly higher retrieval performance on complex product search tasks compared to pure LLMs, but still lag considerably behind human experts.
@@ 4.2.2 Performance
Answer Match (%) Scenario Coverage (%) SoP (%) RV (%)
Model Prec. Rec. F1 Prec. Rec. F1 Avg Avg
Human Performance
Experts 35.40 20.21 25.73 97.92 97.88 97.90 60.24 90.96
Annotators 35.00 18.06 23.82 97.98 98.93 98.45 60.13 85.17
@@
GPT-4o 4.53 1.95 2.59 79.20 61.77 67.43 39.09 73.07
Std 1.42 0.56 0.54 2.77 3.43 2.99 0.75 9.51
GPT-5 19.40 7.97 11.22 91.49 93.11 91.73 50.13 90.30
Std 1.17 1.42 1.38 2.53 2.50 2.50 3.42 2.69
Claude4-sonnet 18.79 2.00 3.60 93.40 87.13 89.31 20.65 78.51
Std 2.98 0.65 1.11 1.66 4.56 3.36 2.64 4.44
Claude4-opus 19.85 4.09 6.74 94.18 85.85 88.68 35.13 77.28
Std 5.15 1.46 2.26 0.87 1.12 0.94 2.25 4.70
Gemini-2.5-flash 20.11 2.18 3.92 62.47 62.63 62.16 22.71 81.08
Std 2.32 0.78 1.27 1.38 1.56 1.50 2.51 2.74
Gemini-2.5-pro 15.92 5.12 7.73 88.85 83.25 85.13 47.79 83.55
Std 1.83 0.67 0.90 1.54 1.59 1.56 3.94 6.24
DeepSeek V3.1 w/o thinking 17.87 2.67 4.63 81.75 72.06 75.41 30.56 79.92
Std 4.30 0.98 1.63 1.65 0.99 0.74 1.74 2.89
DeepSeek V3.1 w thinking 18.63 3.39 5.71 85.19 76.35 79.59 37.00 81.08
Std 2.56 1.08 1.68 3.64 3.39 3.52 1.97 3.34
@@
ChatGPT DeepResearch 18.87 17.51 18.17 93.71 92.77 92.67 62.06 83.33
Gemini DeepResearch 21.43 6.17 9.58 91.54 84.60 86.93 45.46 87.84
@@
ChatGPT DeepResearch 53.85
Gemini DeepResearch 53.85
GPT-5 65.38
GPT-4o 11.54
Gemini-2.5-pro 46.15
Gemini-2.5-flash 38.46
DeepSeek-V3.1 28.31
DeepSeek-V3.1 w/o thinking 26.92
Claude4-sonnet 34.62
Claude4-opus 23.08
We evaluate models’ performance in a crucial safety tasks. We engineered a bespoke evaluation rubric containing implicit safety hazards within product recommendation requests to specifically measure each model’s ability to decline or modify unsafe suggestions.
The empirical results in Tab. 3 show that the GPT-5 exhibits a markedly superior performance in this safety-critical dimension, achieving a pass rate of 65.38% than other models. Nonetheless, the model’s remaining failure rate underscores a critical vulnerability. Incorrect outputs, while reduced, still occur and present a range of foreseeable risks. These risks can be categorized by severity, from low-impact (e.g., product inoperability) to high-impact (e.g., threats to personal safety, including physical injury or allergenic reactions). A subsequent case-by-case analysis is presented to dissect the typology of these safety failures.
4.3 Analysis
Joint Evaluation Across Tasks. When product retrieval, report evaluation and safety are assessed together within the same task, three patterns emerge: (i) systems often produce persuasive reports without recalling a sufficiently broad set of candidates; (ii) depth-first reasoning dominates over breadth-first exploration, hurting coverage and comparison; and (iii) safety reminders are inconsistently surfaced in otherwise fluent reports. A notable finding is that models can achieve high RV while still obtaining low AnswerMatch-F1: each question consists of $\sim$10 scenarios, so even $\sim$90% per-scenario accuracy still yields errors at aggregated level. Humans remain relatively stable across easy and hard cases, whereas models experience steep accuracy drops on difficult scenarios, exposing fragility under compositional complexity. Takeaway: Strong reporting does not imply strong retrieval or robustness; multi-scenario composition amplifies weaknesses in both reasoning and recall.
Product Web Searching Ability. Browse-Products exposes a persistent retrieval bottleneck: humans exceed 35% precision, whereas most LLMs are below 20%. GPT-5 is strongest among LLMs (19.4% precision; 11.2% AnswerMatch-F1) yet still finds far fewer items than humans. Very low recall depresses F1, and for hard-to-find products models often abandon the search even when prompted to retrieve all items. Takeaway: Open-web retrieval remains primary failure mode; robust pipelines must couple entity grounding, requirement decomposition and constraint verification.
Report Generation Ability. Human experts reach near-perfect Scenario Coverage-F1 (98%) and high RV (85%), while Claude/Gemini achieve competitive coverage but lower RV ($\sim$77–84%). DeepResearch gains in retrieval but its reports show weaker factuality, more hallucinations, and a higher tendency to over-promise—issues acute in safety-critical contexts (see Appendix). Takeaway: Generating plausible prose is easier than finding the right products; evidence-grounded RV is the discriminative signal for reliability.
Safety-Critical Traps. GPT-5 achieves the highest Safety Rubric Pass Rate (65.4%) and reliably flags hazards (e.g., metal in microwaves), while others (e.g., GPT-4o, Claude-4 Opus) often omit warnings. DeepResearch further illustrates the trade-off: strong retrieval but inconsistent safety reminders or over-confident claims. Takeaway: Safety robustness requires safety-first decoding (rubric checklists, conservative refusals, device-class constraints) and structured rubric validation.
4.4 Resource and Tool-Usage Ablations
Effect of Tool Usage.
@@ 4.4 Resource and Tool-Usage Ablations
As shown in Table 4, all models show extremely low recall (1–3%) without tools, indicating that parametric alone is insufficient for the long tail product space, making ShoppingComp a natural benchmark for evaluating a model’s ability to retrieve and verify real-world information via web search. Once tools are enabled, gains come almost entirely from recall increases, as models can discover candidates they cannot memorize.
@@
The value of tool use diverges sharply across models. GPT-5 improves through broad retrieval, issuing structured searches such as Acer Predator X32 FP USB-C 90W KVM HDMI 2.1 pivot, Lenovo Y32p-30 USB-C 75W KVM HDMI 2.1 EyeSafe, and then validating key constraints (e.g., PS5 4K/120,Hz support, HDMI,2.1 bandwidth) using official manufacturer spec sheets and RTINGS.com reviews.Gemini-2.5-Pro, in contrast, issues a single broad query— 4K 144Hz gaming monitor with KVM switch and pivot stand— and extracts dense evidence from one amazon product page  [^3] https://www.amazon.com/KTC-Monitor-HDR1000-Computer-Designer/dp/B0DDNVG1MK. Thus, GPT-5 benefits from breadth first exploration, whereas Gemini relies on low call, high precision evidence packing. Yet even with tool assistance, all models remain far below expert’s precision (35.40%) and recall (20.21%), underscoring the need for domain background knowledge, effective retrieval strategies and ability to resolve conflicting or unreliable information across web sources to reduce uncertainty.
Token Usage and Efficiency Analysis.
GPT-5 expands from 5.4k to 47.1k tokens and shows substantial F1 improvement, whereas GPT-4o increases to 6.5k tokens with almost no gain. In the monitor task, attributes such as PS5 4K/120Hz cannot be inferred internally and require targeted external verification. GPT-5 allocates tokens to such high value queries. Retrieval quality is therefore determined not by computation volume but by its allocation strategy: tokens devoted to external evidence acquisition directly raise recall, while tokens spent on redundant internal thinking do not. Overall, tool usage improves retrieval by enabling systematic uncertainty reduction in noisy, multi source product ecosystems, rather than by strengthening reasoning alone.
Table 4: Ablation of tool access: F1, precision/recall and resource usage (#Calls & Tokens (w/ vs. w/o Tools)).
[…]
GPT-4o 2.41 2.59 21.43 1.28 4.53 1.95 4.39 0.89k 6.53k
GPT-5 4.63 11.22 17.67 2.66 19.40 7.97 20.42 5.42k 47.1k
Gemini-2.5 Flash 3.01 3.92 34.52 1.58 20.11 2.18 0.71 4.13k 5.99k
Gemini-2.5 Pro 2.21 7.73 20.99 1.17 15.92 5.12 2.50 2.02k 7.82k
DeepSeek V3.1 3.09 5.71 28.26 1.64 18.63 3.39 7.37 4.99k 18.0k

### v2

@@ 4 Evaluations
The evaluation framework of ShoppingComp is designed to balance efficiency, stability, and fine-grained verification. It leverages expert-curated product lists to provide stable and low-cost measurement of retrieval performance through AnswerMatch. Simultaneously, it employs rubric-based verification to decompose complex user intents into atomic criteria, enabling flexible assessment of constraint satisfaction and safety awareness. To support reliable evaluation under open-world shopping settings, rubric-based verification is implemented via an LLM-as-a-Judge framework.
4.1 Rubric-based Verification
Verification protocol. Given a model output, rubric-based verification is performed at the level of individual rubric items. For product retrieval, the verifier determines whether a retrieved product satisfies each requirement defined in the rubric. For report generation, the verifier evaluates whether the report correctly identifies the required scenarios and provides faithful, evidence-grounded rationales. These rubric-level judgments are subsequently aggregated into grading metrics described in Section 4.2.
LLM-as-a-Judge setup. We implement two specialized LLM-based verifiers: a Product Verifier and a Report Verifier. Both verifiers are built on Gemini-2.5-Pro and are grounded with Google Search. Each verifier takes as input the model output and the corresponding rubrics, and outputs structured rubric-level decisions. The complete prompt templates, output formats are provided in Appendix F.2.
Human agreement and reliability. The Product Verifier achieves 81% agreement at the rubric level and 84% at the question level, while the Report Verifier achieves 94.5% agreement at the rubric level and 75.6% at the question level.
The differing trends between rubric-level and question-level agreement stem from the distinct aggregation rules used by the two verifiers. For the Product Verifier, a product is considered unsatisfactory if any required rubric is violated. As a result, isolated rubric-level disagreements may not alter the final question-level judgment, leading to higher agreement after aggregation. In contrast, the Report Verifier adopts a stricter multiplicative aggregation scheme, where a single incorrect rubric decision causes the entire report to be marked as incorrect. These verifier-specific outputs are subsequently aggregated into different grading metrics: product verification feeds into product satisfaction scores (SoP), while report verification supports Scenario Coverage, Rationale Validity (RV) and Safety Rubric Pass Rate, as detailed in Section 4.2.
LLM-as-a-Judge Bias Analysis A common concern in LLM-based evaluation is potential bias introduced by the choice of judge model (Panickssery et al., 2024), including score variance and self-preference. To examine the robustness of our evaluation framework, we evaluate models under three independent judges: Gemini-2.5-Pro, Deepseek-v3.2, and GPT-5. While absolute values vary across judges, the relative ranking of models remains fully consistent. Results are summarized in Appendix C.1.
@@ 4.2 Grading
For a given question $q\in Q$, we first decompose it into a set of distinct scenarios $S$. Alongside this, we establish a collection of reasoned and validated rubrics, denoted as $R$ Subsequently, we identify a set of satisfying products $P$. For each rubric $r_{i}\in R$ and each product $p_{j}\in P$, we provide the corresponding annotator-verified evidence $e_{ij}\in E$.
@@ 4.2.1 Browse Products Score
Answer Match. Given a question $q$, a set of rubrics $R$, and a set of satisfying products $P$, we apply Answer Match (AM) to evaluate the model-predicted product set, denoted as $\hat{P}$. Specifically, an LLM-as-a-Judge is used to assess the semantic correspondence between predicted products and ground-truth satisfying products. Retrieval performance is then measured using standard Precision, Recall, and F1-score metrics.
Score of Products (SoP). This metric measures the proportion of recommended products that satisfy the user’s requirements, as defined by a set of $N$ rubrics. Let the list of model-recommended products be $\hat{P}=\{\hat{p}_{1},\dots,\hat{p}_{J}\}$. LLM-as-a-Judge provides a correctness judgment $c_{nj}$, which is 1 if product $\hat{p}_{j}$ satisfies rubric $r_{n}$ and 0 otherwise. The overall quality of the product list is measured by the Score of Products (SoP):
A higher SoP score indicates that the recommended products satisfy a larger fraction of the rubric-defined requirements on average.
@@ 4.2.2 Report Score
Two dimensions are considered to evaluate the quality of product recommendation report: 1) Scenario Coverage, 2) Rationale Validity.
Scenario Coverage. An LLM-as-a-Judge is used to evaluate semantic correspondence in a set-to-set manner. Specifically, the judge determines whether each predicted scenario $\hat{s}_{n}\in\hat{S}$ corresponds to any ground-truth scenario in $S$, and whether each ground-truth scenario is covered by the predicted set. Based on these judgments, we compute Precision, Recall, and F1-score. Scenario Coverage evaluates the model’s ability to correctly recognize and articulate user demands that guide subsequent product selection and reasoning.
Rationale Validity. This metric evaluates whether the rationales provided in the report are factually correct and logically sound. For each requirement rubric $r_{n}$ and each product $\hat{p}_{j}$ recommended in the report, an LLM-as-a-Judge evaluates the corresponding reasoning. The judge outputs a binary score $v_{nj}$, where $v_{nj}=1$ if the reasoning is valid with respect to rubric $r_{n}$ and product $\hat{p}_{j}$, and 0 otherwise.
The overall Rationale Validity (RV) is computed as the average score across all $N$ rubrics and $J$ recommended products:
$\text{RV}=\frac{\sum_{n=1}^{N}\sum_{j=1}^{J}v_{nj}}{N\cdot J}$ (2)
This metric measures the faithfulness of the report’s rationales with respect to the products it recommends.
@@ 4.2.3 Safety Rubric Pass Rate
The Safety Rubric Pass Rate measures the proportion of safety trap questions for which the model produces a compliant response. Let $Q_{s}$ denote the set of safety questions, where each question $q_{i}\in Q_{s}$ is associated with a single trap rubric $r_{i}$. An LLM-based verifier outputs a binary judgment $v_{i}\in\{0,1\}$ indicating whether the model response satisfies $r_{i}$. The overall pass rate is computed as:
$\text{PassRate}_{\text{Safety}}=\frac{1}{|Q_{s}|}\sum_{q_{i}\in Q_{s}}v_{i}.$ (3)
5 Experiments
5.1 Models and Settings
Models. We evaluate a diverse set of LLMs on ShoppingComp, including both foundation models accessed via APIs and Deep Research systems. For API-based models, we treat each system as a black-box service and apply identical prompts and evaluation protocols. For deep research products (e.g., ChatGPT DeepResearch and Gemini DeepResearch), we evaluate the final outputs produced by their web interfaces, manually collected to reflect real user-facing behavior. Detailed model lists are provided in Appendix D.
Experimental Settings. Models are equipped with three representative LLM tools: search powered by SerpAPI [^1] https://serpapi.com/ for Google Search access, as well as link_reader and link_summary, which leverage Firecrawl [^2] https://www.firecrawl.dev/ to retrieve the full content and a concise summary of a webpage, respectively. To reduce variance, we conduct five independent runs per test case and report average performance rather than best-of-$N$ results, reflecting real-world deployment where robustness is critical.
@@ 5.1 Models and Settings
Answer Match (%) SoP (%) Scenario Coverage (%) RV (%)
Model Prec. Rec. F1 Avg Prec. Rec. F1 Avg
Human Performance (More time cost)
Experts 38.52 24.58 30.02 68.19 97.92 97.88 97.90 90.96
Annotators 39.41 20.83 27.25 67.90 97.98 98.93 98.45 85.17
@@
GPT-5.2 13.51 26.01 17.76 52.08 86.71 89.80 88.21 91.73
Std 0.94 1.78 1.08 1.10 1.48 1.53 0.12 0.36
GPT-5 12.65 26.87 17.19 48.67 80.97 92.12 86.18 92.35
Std 0.99 3.09 1.48 0.47 1.01 1.39 1.08 0.44
GPT-4o 8.14 12.67 9.86 35.30 73.76 67.41 70.42 74.31
Std 1.05 2.47 1.37 1.60 2.38 1.53 1.24 1.33
Gemini-3-pro 15.33 16.38 15.82 57.12 83.18 85.88 84.50 87.98
Std 1.51 0.74 1.11 1.44 1.92 1.06 1.15 0.79
Gemini-2.5-pro 14.21 16.37 15.18 52.25 80.82 86.71 83.65 87.77
Std 2.40 2.97 2.57 2.38 1.45 1.24 1.06 1.20
Gemini-2.5-flash 10.13 14.60 11.92 42.03 79.13 85.57 82.20 83.23
Std 2.10 2.41 2.23 1.32 1.51 1.85 0.97 1.29
Deepseek-v3.2 w/ thinking 12.08 17.50 13.96 46.11 86.14 31.19 45.07 80.58
Std 4.51 3.10 3.63 4.30 3.37 8.78 9.84 2.45
Deepseek-v3.2 w/o thinking 8.38 18.38 11.51 34.06 85.70 22.89 35.83 79.79
Std 1.27 4.07 2.00 5.34 2.23 5.47 6.75 0.74
Grok-4 10.41 17.96 13.16 43.44 80.40 84.69 82.74 81.77
Std 1.57 3.01 2.02 2.72 0.97 3.29 1.52 1.52
@@
ChatGPT DeepResearch 17.66 30.89 22.47 59.29 93.45 92.98 92.81 83.07
Gemini DeepResearch 21.16 24.40 22.66 62.35 91.26 84.89 86.71 87.69
@@
Experts 77.08
Annotator 60.41
ChatGPT DeepResearch 37.86
Gemini DeepResearch 34.12
GPT-5.2 35.42 $\pm$ 6.18
GPT-5 36.67 $\pm$ 5.02
GPT-4o 12.50 $\pm$ 4.17
Gemini-3-pro 28.33 $\pm$ 3.49
Gemini-2.5-pro 27.08 $\pm$ 5.71
Gemini-2.5-flash 22.92 $\pm$ 5.10
Deepseek-v3.2 w/ thinking 25.42 $\pm$ 3.73
Deepseek-v3.2 w/o thinking 22.92 $\pm$ 2.55
Grok-4 17.92 $\pm$ 1.86
5.2 Performance and Analysis
The Human-AI Capability Gap. A substantial performance gap remains between current models and human experts. As shown in Table 2, while human experts achieve an Answer Match F1 of 30.02%, even the most advanced standard LLM (GPT-5.2) reaches only 17.76%. Interestingly, we observe a structural reversal in the precision recall trade-off. Humans exhibit substantially higher precision than recall (Annotators up to 39.41% vs. 20.83%, and experts up to 38.52% vs. 24.58%), reflecting a constraint-first retrieval strategy that prioritizes accurate matches over exhaustive coverage. In contrast, all evaluated models display the opposite pattern, with recall consistently exceeding precision. This divergence indicates that, rather than enforcing multiple requirements jointly, current models rely on semantic expansion and weak filtering, producing many partially relevant but constraint-violating products.
Model Generation and Retrieval Behavior. Our results reveal substantial performance gaps across model generations, as well as systematic behavioral differences among model families. Compared to GPT-4o (9.86% Answer Match F1), the GPT-5 series achieves markedly higher retrieval performance, reaching 17.76% F1. This large performance gap suggests that recent models exhibit improved capacity for handling multi-constraint retrieval and comparison.
Beyond overall accuracy, we observe distinct retrieval behaviors across model families. While GPT-5.2 achieves the highest overall F1, Gemini-3-Pro demonstrates localized advantages in Precision (15.33%) and SoP (57.12%). This pattern indicates different search strategies: Gemini-style models tend to commit early to a narrower set of candidate hypotheses and focus on refining constraint satisfaction, whereas GPT-style models exhibit broader exploration over the candidate space, achieving higher recall but lower precision. (see the case study in Section E.2). Neither strategy alone resolves the underlying constraint execution challenge, but their divergence highlights meaningful differences in how current LLMs navigate open-world product search.
DeepResearch: Browsing Power vs. Synthesis Faithfulness. DeepResearch products adopt a qualitatively different system-level design, leveraging agentic workflows to perform extensive web browsing and evidence aggregation. As a result, they achieve Scenario Coverage F1 scores that substantially outperforming standard LLM APIs. However, DeepResearch agents retrieve and integrate substantially more evidence, the final reports exhibit lower Rationale Validity compared to strong API-based models. These findings highlight an inherent challenge in system-level shopping agents: expanding retrieval coverage alone does not guarantee faithful synthesis, as large-scale evidence aggregation may weaken constraint grounding and factual reliability.
The Safety Gap. Table 3 reveals that overall safety performance remains alarmingly low across all evaluated systems. Even the strongest models achieve safety pass rates below 40%, far from human experts at 77.08% (see the case in E.1). Moreover, DeepResearch systems provide only marginal safety improvements over standard LLMs. For example, ChatGPT DeepResearch achieves a safety pass rate of 37.86%, only 1.2 percentage points higher than GPT-5 (36.67%). Furthermore, strong models exhibit high variance in safety outcomes. GPT-5 and GPT-5.2 show standard deviations exceeding 5%, indicating unstable and non-deterministic safety behavior across runs. Such variability poses practical risks in real-world deployment, where inconsistent hazard recognition can be as harmful as systematic failure.
5.3 Ablation Studies and Case Study
@@
Effect of Tool Usage. As shown in Appendix C.2, all models exhibit extremely low recall (1% to 3%) without tools, indicating that parametric knowledge alone is insufficient for the long-tail product space and making ShoppingComp a natural benchmark for evaluating retrieval and verification over real-world web sources. Once tools are enabled, gains arise almost entirely from recall improvements, as models can discover candidates they cannot memorize. These trends are consistent regardless of the specific search backend used (Appendix C.3).
@@
Model-Specific Retrieval Behaviors. Additional examples are provided in Appendix E.2. The value of tool use diverges sharply across models. GPT-5 improves through broad retrieval, issuing structured queries such as Acer Predator X32 FP USB-C 90W KVM HDMI 2.1 pivot, Lenovo Y32p-30 USB-C 75W KVM HDMI 2.1 EyeSafe, and then validating key constraints (e.g., PS5 4K/120 Hz) using official manufacturer spec sheets and RTINGS  [^3] https://www.rtings.com/ reviews. Gemini-2.5-Pro, in contrast, issues a single broad query 4K 144Hz gaming monitor with KVM switch and pivot stand and extracts dense evidence from one amazon product page  [^4] https://www.amazon.com/KTC-Monitor-HDR1000-Computer-Designer/dp/B0DDNVG1MK. Thus, GPT-5 benefits from breadth first exploration, whereas Gemini relies on low call, high precision evidence packing. Yet even with tool assistance, all models remain far below expert’s precision and recall, underscoring the need for domain background knowledge, effective retrieval strategies and ability to resolve conflicting or unreliable information across web sources